Divide 44 degrees 8 minutes into two parts such that the no. Of sexagesimal seconds in one part may be equal to the no. Of centesimal seconds in other part.
Please provide the solution.

Grade:11

2 Answers

bharat bajaj IIT Delhi
10 years ago
Say there are x centesimal seconds in one part and x sexagesimal seconds in the other.
Then the total number of degrees would be x / 10000 + x / 3600 = 44 + 8/60.

x = 116823.53 seconds.
Emimilili
6 years ago
a =83621,05263 ".
you have a =first angle in sec sexa and b = lhe other angle in cent (grad)  → (a" and b g.).
a = b (value but Remember b is in g. so b g. = (9/10 ) b ")
but with units you have a +(9/10)b = 44° 8' =((44×60)+8 )X60 = 158 880"  .
 
1f a = b, Then we can also write a + ( 9/10)a = 158  880 → a = 83621, 05263"
so b= 83621, 05263 g. = 83621,05263 × ( 9/10) = 7525 8,94737"
 
You can ckeck.: 
a = 83621, 05263" / (60×60 )=23,22807018 °= 23 ° 13' 41, 05263"
b =83621,05263g, = 7  5258, 94737"/ (60X 60 )= 20° 54' 18,94 737" (ps: b = 25g.80' 89,6686")
 
a+b= 23° 13' 41,05 263" + 20 ° 54' 18, 94737" =  43°67'60''= 44 ° 8 = 83621, 05263" + 75258, 94737" = 158  880"
with same seconds in " OR g, : 83621 , O5263
 
SoRRY for my english, trying to learn, it's why am here